can someone solve for T please
T=6.23*10^8/yz if y=3.8*10^4 and z=2.57*10^3
MY ANSWER but i dont think its right T=6.37927503583862

Answer by stanbon(75887)

T=6.23*10^8/yz if y=3.8*10^4 and z=2.57*10^3
T=6.23E8/[3.8E4*2.57E3]
T=6.23E8/[9.7666E7]
T=6.37927504E0
Cheers,
Stan H.